# History, Development, and Principles of Large Language Models—An Introductory Survey

### v1

@@ 1. INTRODUCTION
Language stands as humanity’s most potent tool, enabling the expression of thoughts and emotions and facilitating communication with others (Pinker, 2023; Turing, 2009). However, machines lack the intrinsic ability to comprehend and communicate in human language, necessitating powerful Artificial Intelligence (AI) algorithms to acquire such capabilities. The ultimate aim of AI research is the creation of machines capable of reading, writing, and conversing like humans (Dwivedi et al., 2023; Zhao et al., 2023; Jin et al., 2023). Natural language processing (NLP) emerges as the AI branch dedicated to realizing this goal. At the core of NLP are language models, tasked with predicting or generating the probability or likelihood of linguistic units (e.g., words, phrases, sentences) based on context. The evolutionary stages of language models unfold chronologically from the initial statistical language models (SLMs) to subsequent neural language models (NLMs), progressing to pre-trained language models (PLMs), and reaching the current state of large language models (LLMs). Specifically, SLMs employ simple probability distributions to model word sequences, while NLMs utilize neural networks to grasp complex patterns and representations of language. PLMs leverage large-scale corpora and self-supervised learning to capture general linguistic knowledge, and LLMs extend PLMs by incorporating massive data, computation, and algorithms, resulting in more expressive, sophisticated, and adaptable language models (Fu et al., 2022;Wei et al., 2022).
In recent years, significant advancements in NLP have been achieved with LLMs, such as the generative pre-trained transformer (GPT) series of models (Radford et al., 2018; Radford et al., 2019; Brown et al., 2020; OpenAI, 2023). Extensively trained on textual data, these models demonstrate the ability to generate human-level texts and perform language-based tasks with exceptional precision. Despite these advancements having greatly benefited various aspects of work and life, LLMs might not be widely understood by practitioners, particularly those without a background in NLP. To provide practitioners with a fundamental understanding of LLMs, this review introduces it across six dimensions: history, development, principles, applications, drawbacks and future directions. This survey distinguishes itself from other reviews on PLMs or LLMs (Zhao et al., 2023; Liu et al., 2023; Han et al., 2021; Shanahan, 2022), which often concentrate on specific aspects using technical language; this survey aims to elucidate LLMs and their principles in a more accessible manner, intending to maximize its full potential.
@@ 2.1 History of LLMs
  • Statistical Language Model: SLMs (Jelinek, 1998; Stolcke, 2002; Rosenfeld, 2000) originated in the 1990s as mathematical models addressing contextually relevant properties of natural language from a probabilistic statistical perspective. The essence of statistical language modeling lies in ascertaining the probability of a sentence occurring within a text. Considering $S$ as the sentence “I am very happy”, $P(\omega_{i})$ signifies the probability of the $i$-th word in the sentence: $\omega_{1}$ as “I”, $\omega_{2}$ as “am”, $\omega_{3}$ as “very”, and $\omega_{4}$ as “happy”. Now, the objective is to ascertain the likelihood of $S$ appearing in the text, denoted as $P(S)$:
@@
  • Neural Language Models: NLMs (Bengio et al., 2000; Mikolov et al., 2010; Kombrink et al., 2011) leverage neural networks to predict the probabilities of subsequent words within sequences. They effectively handle longer sequences and mitigate the limitations associated with small $n$ in SLMs. Before delving into neural networks, let’s grasp the concept of word vectors. Humans effortlessly comprehend word meanings. For instance, “cat” and “dog” share closer semantic connections than “cat” and “tree” since they both represent animals. But how does a computer accomplish this? Computers operate using binary code — 0s and 1s — so human language needs translation into binary. Word vectors can accomplish this, which are numerical representations of human language, where each word corresponds to a distinct vector. These vectors usually possess fixed dimensions and can simulate word relationships through the angles between them. An intriguing example is the angle between the word vectors for “cat” and “dog” which is smaller than the angle between the word vector for “cat” and “tree”. Another illustrative example is “$M\ essi-Argentina+P\ ortugal=Cristiano\ Ronaldo$”. Word2Vec (Mikolov et al., 2013) is a widely recognized tool for computing word vectors. Its function involves converting words into dense numerical representations, making words with similar semantics closer together in the vector space.
@@
  • Pre-trained Language Model: PLMs undergo initial training using an extensive volume of unlabeled text, enabling them to grasp fundamental language structures such as vocabulary, syntax, semantics, and logic — a phase termed pre-training. Subsequently, this comprehensive language model can be applied to various NLP tasks like machine translation, text summarization, and question-answering systems. To optimize its performance, models need to be trained a second time on a smaller dataset customized for a specific downstream task — a phase known as fine-tuning. This is the “pre-training and fine-tuning” learning paradigm. We can use a visual example to understand the “pre-training and fine-tuning”, as follows: in martial arts novels, a person who wants to become a martial arts master needs to have a solid foundation of internal martial arts, which can be acquired by training on a large variety of techniques. Then, the person can learn a specific skill, such as a sword or a palm strike, and master it quickly and effectively by applying internal martial arts principles. A large number of studies on PLMs have been built on this paradigm, which introduces different architectures (Devlin et al., 2018; Fedus et al., 2022) (e.g., GPT-2 (Radford et al., 2019) and Bert (Devlin et al., 2018)).
  • Large Language Model: LLMs are trained on massive text corpora with tens of billions (or more) of parameters, such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), PaLM (Chowdhery et al., 2022), and LLaMA (Touvron et al., 2023). The goal of LLMs is to enable machines to understand human commands and adherence to human values. Their hallmark lies in the consolidation of two stages: initial pre-training on a vast general-purpose corpus followed by alignment with human values, rather than transitioning to a different domain. LLMs exhibit remarkable adaptability compared to PLMs, transitioning from specialized to general-purpose models. The substantial increase in model size, dataset volume, and computational prowess has resulted in significant enhancements across various tasks and unveiled remarkable capabilities absent in smaller models. For example, GPT-3 has the capability to leverage contextual information, a functionality that GPT-2 (Radford et al., 2019) lacks. This means that when GPT-3 is prompted with task-related examples, it utilizes them to enhance its problem-solving capabilities. The number of parameters for LLMs typically exceeds a hundred billion, and the training data is usually in the range of a few hundred GB to a few TB. A concrete example is that there are multiple versions of the GPT-2 model, with the largest version having 1.5 billion parameters and using 40GB of text data for training. In contrast, the largest version of the GPT-3 model has 175 billion parameters and uses 570GB of text data for training. This example illustrates the significant discrepancy between LLMs and PLMs concerning parameter count and training data volume.
@@ 2.2 Factors Propelling Large Language Models
  • Data Diversity: data diversity has emerged as a crucial catalyst for the development of LLMs. Recent years have witnessed a surge in the availability and diversity of extensive internet-based data sources, furnishing these models with training data containing rich linguistic and worldly insights. The quality and origins of this data significantly influence the efficacy and capacity of LLMs. Earlier studies trained language models on a single domain of text, such as news articles (Jozefowicz et al., 2016), Wikipedia (Merity et al., 2016), or fiction books (Kiros et al., 2015). However, the proliferation of vast and heterogeneous text corpora across the web—including blogs, social media, forums, reviews, among others—has spurred researchers to pivot towards training models on multi-domain texts. This shift enhances the model’s aptitude for generalization and multitasking. An exemplary instance is Meta AI’s open-source large language model, LLaMA (Touvron et al., 2023), trained on publicly accessible datasets like CommonCrawl [^1] https://commoncrawl.org/get-started, C4 (Raffel et al., 2020), Github [^2] https://github.com/, Wikipedia [^3] https://huggingface.co/datasets/wikipedia, Books (Gao et al., 2020), ArXiv [^4] https://arxiv.org/, and StackExchange [^5] https://stackexchange.com/.
@@
  • Computational Advancement: LLMs are neural network models with a large number of parameters. With the progression of deep learning (DL) technology in recent years, the scale and effectiveness of LLMs have notably escalated. Simultaneously, they present serious challenges due to their substantial computational demands (Kaplan et al., 2020). Computational power signifies a computer system’s capability to execute computational tasks, often quantified in terms of floating-point operations per second (Kaplan et al., 2020). According to OpenAI calculations [^6] https://openai.com/research/ai-and-compute, since 2012, the amount of computation used for global AI training has shown exponential growth, doubling every 3.43 months on average, and now the amount of computation has expanded 300,000 times, far outpacing the rate of computational power growth. Fortunately, hardware innovators like Nvidia continually invent specialized devices such as GPUs, TPUs, and NPUs with amplified computational prowess, LLMs can be trained quickly. For instance, the launch of GPT-3 by OpenAI in May 2020 (Brown et al., 2020) underscores this reliance on robust hardware. Training GPT-3 on a single NVIDIA Tesla V100 GPU would require an estimated 355 years. Conversely, leveraging 1024×A100 GPUs, researchers estimated the potential to train GPT-3 in as little as 34 days. This exemplifies the indispensability of potent hardware in facilitating the development and training of LLMs.
@@
  • Algorithmic Innovation: algorithmic innovation stands as the pivotal driving force behind LLMs, shaping their structure and functionality. From the initial rule- and statistics-based approach, to the later DL-based approach, the algorithms of language models have continuously evolved and improved. Presently, all LLMs are based on the transformer architecture (Vaswani et al., 2017), which employs a self-attention mechanism (see discussion in Section 3). This architecture offers distinct advantages over traditional recurrent (Mikolov et al., 2010) and convolutional neural networks (LeCun et al., 1998), enabling fully parallel computation and adeptly capturing long-distance dependencies. Moreover, there exist numerous iterations and enhancements of the transformer, such as Transformer-XL (Dai et al., 2019), XLNet (Yang et al., 2019), ALBERT (Lan et al., 2019), each targeting specific aspects for optimization. These variations aim to enhance the attention mechanism, refine pre-training objectives, and curtail parameter counts, thereby advancing the transformer architecture in diverse dimensions.
@@ 3. PRINCIPLES OF LARGE LANGUAGE MODELS
This section strives to clarify the fundamental principles of LLMs, intending to provide general practitioners with a more profound understanding of the operational mechanisms employed by these models. Fundamental to LLMs is the utilization of deep neural networks to grasp the statistical patterns and semantic nuances of language, enabling the understanding and generation of language. The GPT series, a notable example among these models, incorporates the transformer architecture (Vaswani et al., 2017) and utilizes autoregression to predict consecutive words. The following discussion will concentrate on the GPT-3 (Brown et al., 2020) as an exemplar, providing insights into multiple facets that elucidate the principles of LLMs, as illustrated in Figure 3.
@@
    where $pos$ is the position of the current word within the sentence (e.g. in the sentence “A survey of Large Language Model”, $pos_{A}$=0, $pos_{Large}$=3), $d_{model}$ is the length of the embedding vector 12288 and $i$ is the dimension. However, why delve into such intricate computations for positional encodings? Let’s explore a more straightforward approach to conceptualize it. Initially, considering a text of length $T$, the most basic approach of the positional encoding is counting. This method utilizes $[0,1,2...T-1]$ as positional encodings for each word in the text. For instance, the positional encodings for the $2$-th word would be $[1,1,...,1]$, aligning in length with the embedding vector. However, this encoding strategy presents a notable drawback: within longer sentences, subsequent words hold considerably larger positional encoding values compared to the positional encodings of the initial words. As a result, this disparity inevitably introduces numerical biases in the merged features when combined with word embeddings. To address this issue, normalization becomes a viable consideration. The most straightforward normalization method involves dividing directly by $T$, resulting in positional encodings such as $[0,\frac{1}{T},\ldots,\frac{T-1}{T}]$ (e.g. the 2nd word’s positional encodings=$[\frac{1}{T},\frac{1}{T},\ldots,\frac{1}{T}]$). While this method restricts all positional encodings to the [0,1] range, it introduces a significant inconsistency: the positional values are contingent on the text’s length, causing the positional distinctions between adjacent words in shorter texts to differ from those in longer texts. As illustrated in Figure 4, the positional encodings discrepancy between adjacent word positions in sentence ① is 0.5, while in sentence ②, neighboring word positions are coded with a difference of 0.125. Consequently, neither of these methods proves to be satisfactory. To address this issue, Google devised an alternative approach involving trigonometric functions (Vaswani et al., 2017). As a result, a positional encoding matrix denoted as $X_{PositionalEncoding}\in\mathbb{R}^{2048\times 12288}$.
@@
    Masked Self-Attention: Masked self-attention is a modified version of self-attention designed to prevent the model from accessing future words during the prediction of the next word by incorporating a masking technique. For example, with the corpus “abcde”, we can train GPT to predict the next word four times by feeding in “a”, “ab”, “abc” and “abcd” separately. However, this involves segmenting and inputting the sentence four times. The mask technique trick streamlines this process, enabling the model to receive the entire sentence at once and generate an output for each word. The challenge arises from the model peeking at subsequent words, which necessitates the use of masking. When predicting the next word of “a” GPT will only see “a”, and when predicting the next word of “ab”, it will only see “ab”. It’s essential to differentiate between self-attention (utilized in BERT (Devlin et al., 2018)) and masked self-attention (utilized in GPT-3 (Brown et al., 2020)). Figure 5 provides a visual illustration.
@@
  • Add & Norm: The add & norm modules are incorporated into each transformer layer on the GPT model architecture, where “add” represents the skip connection (He et al., 2016) and “norm” represents layer normalization (Ba et al., 2016). The skip connection helps alleviate the issue of gradient vanishing in deep models, enabling the transformer layer to be stacked to a considerable depth. Additionally, layer normalization serves as a regularization strategy to prevent network overfitting. It calculates the mean and variance on each sample independently, without considering other data, thereby mitigating the impact of varying batch sizes.

### v2

@@ 1. INTRODUCTION
Language stands as humanity’s most potent tool, enabling the expression of thoughts and emotions and facilitating communication with others [pinker2003language, turing2009computing]. However, machines lack the intrinsic ability to comprehend and communicate in human language, necessitating powerful Artificial Intelligence (AI) algorithms to acquire such capabilities. The ultimate aim of AI research is the creation of machines capable of reading, writing, and conversing like humans [dwivedi2023so, zhao2023survey, jin2023rethinking]. Natural language processing (NLP) emerges as the AI branch dedicated to realizing this goal. At the core of NLP are language models, tasked with predicting or generating the probability or likelihood of linguistic units (e.g., words, phrases, sentences) based on context. The evolutionary stages of language models unfold chronologically from the initial statistical language models (SLMs) to subsequent neural language models (NLMs), progressing to pre-trained language models (PLMs), and reaching the current state of large language models (LLMs). Specifically, SLMs employ simple probability distributions to model word sequences, while NLMs utilize neural networks to grasp complex patterns and representations of language. PLMs leverage large-scale corpora and self-supervised learning to capture general linguistic knowledge, and LLMs extend PLMs by incorporating massive data, computation, and algorithms, resulting in more expressive, sophisticated, and adaptable language models [fu2022does, wei2022emergent].
In recent years, significant advancements in NLP have been achieved with LLMs, such as the generative pre-trained transformer (GPT) series of models [radford2018improving, radford2019language, brown2020language, achiam2023gpt]. Extensively trained on textual data, these models demonstrate the ability to generate human-level texts and perform language-based tasks with exceptional precision. Despite these advancements having greatly benefited various aspects of work and life, LLMs might not be widely understood by practitioners, particularly those without a background in NLP. To provide practitioners with a fundamental understanding of LLMs, this review introduces it across six dimensions: history, development, principles, applications, drawbacks and future directions. This survey distinguishes itself from other reviews on PLMs or LLMs [zhao2023survey, liu2023pre, han2021pre, shanahan2024talking], which often concentrate on specific aspects using technical language; this survey aims to elucidate LLMs and their principles in a more accessible manner, intending to maximize its full potential.
@@ 2.1 History of LLMs
  • Statistical Language Model: SLMs [jelinek1998statistical, stolcke2002srilm, rosenfeld2000two] originated in the 1990s as mathematical models addressing contextually relevant properties of natural language from a probabilistic statistical perspective. The essence of statistical language modeling lies in ascertaining the probability of a sentence occurring within a text. Considering $S$ as the sentence “I am very happy”, $P(\omega_{i})$ signifies the probability of the $i$-th word in the sentence: $\omega_{1}$ as “I”, $\omega_{2}$ as “am”, $\omega_{3}$ as “very”, and $\omega_{4}$ as “happy”. Now, the objective is to ascertain the likelihood of $S$ appearing in the text, denoted as $P(S)$:
@@
  • Neural Language Models: NLMs [bengio2000neural, mikolov2010recurrent, kombrink2011recurrent] leverage neural networks to predict the probabilities of subsequent words within sequences. They effectively handle longer sequences and mitigate the limitations associated with small $n$ in SLMs. Before delving into neural networks, let’s grasp the concept of word vectors. Humans effortlessly comprehend word meanings. For instance, “cat” and “dog” share closer semantic connections than “cat” and “tree” since they both represent animals. But how does a computer accomplish this? Computers operate using binary code — 0s and 1s — so human language needs translation into binary. Word vectors can accomplish this, which are numerical representations of human language, where each word corresponds to a distinct vector. These vectors usually possess fixed dimensions and can simulate word relationships through the angles between them. An intriguing example is the angle between the word vectors for “cat” and “dog” which is smaller than the angle between the word vector for “cat” and “tree”. Another illustrative example is “$M\ essi-Argentina+P\ ortugal=Cristiano\ Ronaldo$”. Word2Vec [mikolov2013distributed] is a widely recognized tool for computing word vectors. Its function involves converting words into dense numerical representations, making words with similar semantics closer together in the vector space.
@@
  • Pre-trained Language Model: PLMs undergo initial training using an extensive volume of unlabeled text, enabling them to grasp fundamental language structures such as vocabulary, syntax, semantics, and logic — a phase termed pre-training. Subsequently, this comprehensive language model can be applied to various NLP tasks like machine translation, text summarization, and question-answering systems. To optimize its performance, models need to be trained a second time on a smaller dataset customized for a specific downstream task — a phase known as fine-tuning. This is the “pre-training and fine-tuning” learning paradigm. We can use a visual example to understand the “pre-training and fine-tuning”, as follows: in martial arts novels, a person who wants to become a martial arts master needs to have a solid foundation of internal martial arts, which can be acquired by training on a large variety of techniques. Then, the person can learn a specific skill, such as a sword or a palm strike, and master it quickly and effectively by applying internal martial arts principles. A large number of studies on PLMs have been built on this paradigm, which introduces different architectures [devlin2018bert, fedus2022switch] (e.g., GPT-2 [radford2019language] and Bert [devlin2018bert]).
  • Large Language Model: LLMs are trained on massive text corpora with tens of billions (or more) of parameters, such as GPT-3 [brown2020language], GPT-4 [achiam2023gpt], PaLM [chowdhery2023palm], and LLaMA [touvron2023llama]. The goal of LLMs is to enable machines to understand human commands and adherence to human values. Their hallmark lies in the consolidation of two stages: initial pre-training on a vast general-purpose corpus followed by alignment with human values, rather than transitioning to a different domain. LLMs exhibit remarkable adaptability compared to PLMs, transitioning from specialized to general-purpose models. The substantial increase in model size, dataset volume, and computational prowess has resulted in significant enhancements across various tasks and unveiled remarkable capabilities absent in smaller models. For example, GPT-3 has the capability to leverage contextual information, a functionality that GPT-2 [radford2019language] lacks. This means that when GPT-3 is prompted with task-related examples, it utilizes them to enhance its problem-solving capabilities. The number of parameters for LLMs typically exceeds a hundred billion, and the training data is usually in the range of a few hundred GB to a few TB. A concrete example is that there are multiple versions of the GPT-2 model, with the largest version having 1.5 billion parameters and using 40GB of text data for training. In contrast, the largest version of the GPT-3 model has 175 billion parameters and uses 570GB of text data for training. This example illustrates the significant discrepancy between LLMs and PLMs concerning parameter count and training data volume.
@@ 2.2 Factors Propelling Large Language Models
  • Data Diversity: data diversity has emerged as a crucial catalyst for the development of LLMs. Recent years have witnessed a surge in the availability and diversity of extensive internet-based data sources, furnishing these models with training data containing rich linguistic and worldly insights. The quality and origins of this data significantly influence the efficacy and capacity of LLMs. Earlier studies trained language models on a single domain of text, such as news articles [jozefowicz2016exploring], Wikipedia [merity2016pointer], or fiction books [kiros2015skip]. However, the proliferation of vast and heterogeneous text corpora across the web—including blogs, social media, forums, reviews, among others—has spurred researchers to pivot towards training models on multi-domain texts. This shift enhances the model’s aptitude for generalization and multitasking. An exemplary instance is Meta AI’s open-source large language model, LLaMA [touvron2023llama], trained on publicly accessible datasets like CommonCrawl [^1] https://commoncrawl.org/get-started, C4 [raffel2020exploring], Github [^2] https://github.com/, Wikipedia [^3] https://huggingface.co/datasets/wikipedia, Books [gao2020pile], ArXiv [^4] https://arxiv.org/, and StackExchange [^5] https://stackexchange.com/.
@@
  • Computational Advancement: LLMs are neural network models with a large number of parameters. With the progression of deep learning (DL) technology in recent years, the scale and effectiveness of LLMs have notably escalated. Simultaneously, they present serious challenges due to their substantial computational demands [kaplan2020scaling]. Computational power signifies a computer system’s capability to execute computational tasks, often quantified in terms of floating-point operations per second. According to OpenAI calculations [^6] https://openai.com/research/ai-and-compute, since 2012, the amount of computation used for global AI training has shown exponential growth, doubling every 3.43 months on average, and now the amount of computation has expanded 300,000 times, far outpacing the rate of computational power growth. Fortunately, hardware innovators like Nvidia continually invent specialized devices such as GPUs, TPUs, and NPUs with amplified computational prowess, LLMs can be trained quickly. For instance, the launch of GPT-3 by OpenAI in May 2020 [brown2020language] underscores this reliance on robust hardware. Training GPT-3 on a single NVIDIA Tesla V100 GPU would require an estimated 355 years. Conversely, leveraging 1024×A100 GPUs, researchers estimated the potential to train GPT-3 in as little as 34 days. This exemplifies the indispensability of potent hardware in facilitating the development and training of LLMs.
@@
  • Algorithmic Innovation: algorithmic innovation stands as the pivotal driving force behind LLMs, shaping their structure and functionality. From the initial rule- and statistics-based approach, to the later DL-based approach, the algorithms of language models have continuously evolved and improved. Presently, all LLMs are based on the transformer architecture [vaswani2017attention], which employs a self-attention mechanism (see discussion in Section 3). This architecture offers distinct advantages over traditional recurrent [mikolov2010recurrent] and convolutional neural networks [lecun1998gradient], enabling fully parallel computation and adeptly capturing long-distance dependencies. Moreover, there exist numerous iterations and enhancements of the transformer, such as Transformer-XL [cui2023chatlaw], XLNet [yang2019xlnet], ALBERT [lan2019albert], each targeting specific aspects for optimization. These variations aim to enhance the attention mechanism, refine pre-training objectives, and curtail parameter counts, thereby advancing the transformer architecture in diverse dimensions.
@@ 3. PRINCIPLES OF LARGE LANGUAGE MODELS
This section strives to clarify the fundamental principles of LLMs, intending to provide general practitioners with a more profound understanding of the operational mechanisms employed by these models. Fundamental to LLMs is the utilization of deep neural networks to grasp the statistical patterns and semantic nuances of language, enabling the understanding and generation of language. The GPT series, a notable example among these models, incorporates the transformer architecture [vaswani2017attention] and utilizes autoregression to predict consecutive words. The following discussion will concentrate on the GPT-3 [brown2020language] as an exemplar, providing insights into multiple facets that elucidate the principles of LLMs, as illustrated in Figure 3.
@@
    where $pos$ is the position of the current word within the sentence (e.g. in the sentence “A survey of Large Language Model”, $pos_{A}$=0, $pos_{Large}$=3), $d_{model}$ is the length of the embedding vector 12288 and $i$ is the dimension. However, why delve into such intricate computations for positional encodings? Let’s explore a more straightforward approach to conceptualize it. Initially, considering a text of length $T$, the most basic approach of the positional encoding is counting. This method utilizes $[0,1,2...T-1]$ as positional encodings for each word in the text. For instance, the positional encodings for the $2$-th word would be $[1,1,...,1]$, aligning in length with the embedding vector. However, this encoding strategy presents a notable drawback: within longer sentences, subsequent words hold considerably larger positional encoding values compared to the positional encodings of the initial words. As a result, this disparity inevitably introduces numerical biases in the merged features when combined with word embeddings. To address this issue, normalization becomes a viable consideration. The most straightforward normalization method involves dividing directly by $T$, resulting in positional encodings such as $[0,\frac{1}{T},\ldots,\frac{T-1}{T}]$ (e.g. the 2nd word’s positional encodings=$[\frac{1}{T},\frac{1}{T},\ldots,\frac{1}{T}]$). While this method restricts all positional encodings to the [0,1] range, it introduces a significant inconsistency: the positional values are contingent on the text’s length, causing the positional distinctions between adjacent words in shorter texts to differ from those in longer texts. As illustrated in Figure 4, the positional encodings discrepancy between adjacent word positions in sentence ① is 0.5, while in sentence ②, neighboring word positions are coded with a difference of 0.125. Consequently, neither of these methods proves to be satisfactory. To address this issue, Google devised an alternative approach involving trigonometric functions [vaswani2017attention]. As a result, a positional encoding matrix denoted as $X_{PositionalEncoding}\in\mathbb{R}^{2048\times 12288}$.
@@
    Masked Self-Attention: Masked self-attention is a modified version of self-attention designed to prevent the model from accessing future words during the prediction of the next word by incorporating a masking technique. For example, with the corpus “abcde”, we can train GPT to predict the next word four times by feeding in “a”, “ab”, “abc” and “abcd” separately. However, this involves segmenting and inputting the sentence four times. The mask technique trick streamlines this process, enabling the model to receive the entire sentence at once and generate an output for each word. The challenge arises from the model peeking at subsequent words, which necessitates the use of masking. When predicting the next word of “a” GPT will only see “a”, and when predicting the next word of “ab”, it will only see “ab”. It’s essential to differentiate between self-attention (utilized in BERT [devlin2018bert]) and masked self-attention (utilized in GPT-3 [brown2020language]). Figure 5 provides a visual illustration.
@@
  • Add & Norm: The add & norm modules are incorporated into each transformer layer on the GPT model architecture, where “add” represents the skip connection [kaiming2016deep] and “norm” represents layer normalization [ba2016layer]. The skip connection helps alleviate the issue of gradient vanishing in deep models, enabling the transformer layer to be stacked to a considerable depth. Additionally, layer normalization serves as a regularization strategy to prevent network overfitting. It calculates the mean and variance on each sample independently, without considering other data, thereby mitigating the impact of varying batch sizes.